OpenStudy (anonymous):

Solve for w 1/2w + 2 =4 A. w = 12 B. w = 4 C. w = 2 D. w = 3

OpenStudy (bahrom7893):

multiply everything by 2

OpenStudy (bahrom7893):

what do u get?

OpenStudy (anonymous):

what do you mean by everything?

OpenStudy (bahrom7893):

every term: \[2*\frac{1}{2}w+2*2=2*4\]

OpenStudy (bahrom7893):

simplify that

OpenStudy (bahrom7893):

Come on it's just some multiplication.

OpenStudy (anonymous):

so i just do 1/2 * 2, 2*2 & 4*2 ??

OpenStudy (bahrom7893):

yes what do those numbers equal to?

OpenStudy (bahrom7893):

yahya please don't give away the answer.

OpenStudy (anonymous):

get free from w by multipling equation by w

OpenStudy (bahrom7893):

You can't get rid of w. Do as I said, @renae_

OpenStudy (anonymous):

1/2 * 2 = 1 2*2 = 4 4*2 = 8 ??

OpenStudy (bahrom7893):

Yes that's what I wanted, but I wanted you to write it this way: \[2*\frac{1}{2}*w + 2*2 = 2*4\]\[1*w+4=8\]\[w+4=8\]Now subtract 4 from both sides. What do you get?

OpenStudy (anonymous):

1w + 4 =8 where do i subtract from?

OpenStudy (bahrom7893):

Here: w+4 - 4 = 8 - 4 Basically an equation is like a scale. Whatever you do to one side, you need to do to the other, to keep the scales balanced. So if you add 5 to one side, you have to add 5 to the other side as well. If you have to subtract 4 from one side, you have to subtract 4 from the other side as well.

OpenStudy (anonymous):

w + 0 = 4? so B?

OpenStudy (bahrom7893):

yes